OpenStudy (vijay222ms):

Near the earth surface PE = mgh and PE = GMm/r is equal .Give an example and prove

OpenStudy (vijay222ms):

@sidsiddhartha

OpenStudy (vijay222ms):

@aryandecoolest

OpenStudy (vijay222ms):

@Chineseboy15 @♪Chibiterasu @tHe_FiZiCx99

OpenStudy (vijay222ms):

@Abhisar @abb0t @jhonyy9 @Preetha @esshotwired @Gabylovesyou

OpenStudy (abb0t):

@wyattp17

OpenStudy (vijay222ms):

@texaschic101

OpenStudy (schrodingers_cat):

Recall that g = GM/Re^2 at a distance from earth. Also U(Re) = -gmM/Re. Subbing these equations into each other you get U(Re) = -mgRe. For very small heights h above the earths surface where h<<Re U(Re + h) = -gmM/(Re + h) is approximately equal to -mg(Re +h). So, the change in potential energy of a falling mass would be \[\Delta U = Uf - Ui\] which is U = (-mgRe - (-mg(Re +h)) U = mgh Hope this helps :)